How many quarts of pure antifreeze must be added to 6 quarts of a 20% antifreeze solution to obtain a 40% antifreeze solution?
Found 2 solutions by Boreal, macston:
Answer by Boreal(15235)

pure antifreeze=x qts.
(6+x)*0.40=6*0.20+x
2.40+0.40x=1.2+x
1.20=0.6x
x=2 qts.
What you want to equalize is the amount of pure antifreeze in both sides of the equation.
you start with 1.2 (6*0.2) qts.
you are adding x qts to 6.
you are ending up with (6+x) qts of 40% or 0.4
8 qts of 0.4=3.2 qts of pure
You start with 6 qts. and end up with 8.

Answer by macston(5194)

.
x=quarts of pure antifreeze to add
.
0.20(6)+x=0.40(6+x)
1.2+x=2.4+0.40x . Subtract (0.40x) from each side.
1.2+0.60x=2.4 . Subtract 1.2 from each side.
0.60x=1.2 . Divide each side by (0.60)
x=2
ANSWER: You should add 2 quarts of pure antifreeze.
.